# Pareto Envelope Augmented with Reinforcement Learning: Multi-objective reinforcement learning-based approach for

### 19 Dec 2023 v2

@@ 5.2.1 Unconstrained Benchmark problems
The details of the geometry, the decision space $\Gamma$, and the feasible space $\mathbb{O}\subset\Gamma$ are given in [1, 2]. As mentioned in Section 4.2.2, we decided to minimize the peaking $F_{\Delta h}$, maximize the $L_{C}$, and minimize the average enrichment in the core. The first test will utilize the first two objectives, while the second test will consider all three. It will be intriguing to observe how the envelope coverage changes as we are adding a third type of trade-off surface. There, we ran the RL agents (similar hyper-parameters for RL as in [1, 2]) from scratch up to 50,000 and 100,000 objective function evaluations for the Two- and Three-obj cases, respectively, while using 32 cores (i.e. 32 agents) on an Intel Xeon Gold 5220R clocked at 2.20 GHz. For cosine similarity as uniformity for PEARL-e, the option normalized_obj is set to False. Similarly, we ran the NSGA-III on 32 cores, which results in 32 individuals in a population. For the GA portion of the NSGA-III algorithms, we used the Evolutionary ($\mu$,$\lambda$) Strategy (ES) [51]. We use $\lambda$_ equal to 32, $\mu$ equal to $\lambda$_,mutpb to 0.3, and cxpb to 0.65. NSGA-II and NSGA-III will be available in a further release of NEORL [52] for reproducibility. Since PEARL will propose more solutions from the different buffers, we will provide the pareto front of all the solutions ever found from NSGA-III rather than the last population (i.e., 32 of them) for fairness. The individual results for the Two- and Three-obj cases are given in Figure 1 on the top and bottom, respectively (grouped in four figures each).
@@
In the Two-obj case, a trade-off between the $L_{C}$ and $F_{\Delta h}$ becomes apparent. Furthermore, when peaking factors are extremely low, the $L_{C}$ can undergo significant variations, even though it is feasible to achieve a comparable $L_{C}$ level for different values of $F_{\Delta h}$, ranging from approximately 1.55 to 2.2. This situation could potentially lead to substantial economic losses if the constraint on $F_{\Delta h}$ is overly stringent. The PEARL-e and PEARL-$\epsilon$ approaches fail to find low $F_{\Delta h}$ and the latter preferentially generated higher $L_{C}$ (or lower $-L_{C}$), while the former is stuck in a low quality Pareto front. The figure alone does not provide a clear indication of which approach (PEARL-NdS (crowding), or NSGA-III) is better. However, the high HV for the PEARL-NdS (crowding) in Table 8 suggests that the latter approach could provide better volume coverage, but also surpasses the NSGA-based approach on all the other indicators. Moreover, if we only look at the PEARL-NdS (crowding) versus NSGA-III (in the figure and $I_{C}$ in Table 8), we see that PEARL will propose more Pareto-optimal points. The use of multiple metrics helps therefore to assess the performance of multi-objective approaches and demonstrate PEARL-NdS (crowding) superiority.
@@
More statistical weights was be given to $F_{\Delta h}$. The five best Pareto-optimal points are given by the three optimum from NSGA-III namely $[1.43,543.2]$, $[1.44,546]$, $[1.45,548.6]$ and two from the PEARL-NdS (crowding)-based approach $[1.456,553.5]$, and $[1.457,553.9]$. Among the PEARL algorithms, this provides an additional reason to consider PEARL-NdS (crowding) with CL as potentially superior. Furthermore, the best candidate has a very high $L_{C}$ compared to the limit of 500 in [1, 2]. However, approaching the limit of 1.45 by turning peaking into a constraint rather than an objective could result in lower $L_{C}$ but also discard the two best PEARL-NdS (crowding) cases mentioned above. Nevertheless, in Section 5.2.2 we will investigate the impact of introducing constraints to guide the search in PEARL for this set of problems and illustrate its potential to still yield high-quality feasible solutions.
Figure 1: Two-obj (upper group) and Three-obj (lower group) solutions generated for all the algorithms utilized. The pareto front is a line for the Two-objective case, while it is drawn with scattered points for the Three-objective one.
@@
For the Three-obj case, most techniques found a Pareto-optimal point (C_metric $\neq 0.0$) except PEARL-e (cos). Nevertheless again, the PEARL-NdS (crowding) approach performs better in terms of HV as seen in Table 9. The high HV for the PEARL-NdS (crowding) technique is also illustrated by Figure 2. There, we observe a significant variance in the objective space explored when using the latter method, while PEARL-$\epsilon$ exacerbate a very low one, resulting in solutions that are concentrated around relatively high $L_{C}$, intermediate $F_{\Delta h}$ and average enrichment. The PEARL-$\epsilon$-based approach outperforms the other algorithms in terms of GD, $I_{\epsilon}^{+}$, $I_{C}$ and C_metric but exacerbates a very poor HV. The latter has found lots of high-quality points (high $I_{C}$ equal to 1755 which is the highest) but these points are highly localized, as corroborated by the high IGD. This underscores the importance of employing multiple multi-objective metrics when analyzing the performance of the algorithms.
@@
Table 8: Different FOMs for the Two-obj case. The nadir is $[2.653,481.2]$ and the utopia is $[1.43,558.0]$. They are obtained from cases generated over the 4 optimization runs. All the indicators are calculated only with the Pareto equivalent points. The C_metric is equal to 0 when no solution belongs to the Pareto front. $I_{C}$ is the number of Pareto-optimal points for the algorithm. The best cases are highlighted in red.
FOMs PEARL-e (cos) PEARL-$\epsilon$ PEARL NdS (crowding) NSGA-III
HV 69.805 82.337 92.054 87.786
GD 4.859 0.920 0.0529 0.739
IGD 12.236 3.810 0.332 1.981
$I_{\epsilon}^{+}$ 9.300 0.400 0.0170 0.0900
$I_{C}$ 0 0 16 3
C_metric 0.0 0.0 0.941 0.375
@@
Table 9: Different FOMs for the Three-obj case. The nadir is $[3.606,4.75,451.5]$ and the utopia $[1.43,3.911,553.4]$. They are obtained from cases generated over the 4 optimization runs. All the indicators are calculated only with the Pareto equivalent points. The C_metric is equal to 0 when no solution belongs to the Pareto front. $I_{C}$ is the number of Pareto-optimal points for the algorithm. The best cases are highlighted in red.
FOMs PEARL-e (cos) PEARL-$\epsilon$ PEARL NdS (crowding) NSGA-III
HV 84.883 75.703 98.693 97.322
GD 0.130 0.00329 0.0288 0.0364
IGD 0.626 7.229 0.138 0.551
$I_{\epsilon}^{+}$ 0.178 0.0699 0.122 0.126
$I_{C}$ 0 1755 1155 302
C_metric 0.0 0.962 0.826 0.716
Figure 2: Evolution of the target FOMs for the Three-objective problem over 100 generations for PEARL. We added vertical grid lines to help with visibility and interpretation of the results.
@@
Additionally, it’s noteworthy that in both optimization scenarios, PEARL-e under-performs when compared to the SO-based approaches, even though it performed well in the classicla benchmark cases. This difference in performance could be attributed to the arbitrary choice of $\alpha=[1.0,1.0,1.0]$. Given that we didn’t possess prior knowledge of the Pareto front’s shape, we opted for a uniform generation of rays. It is conceivable that a designer could iteratively construct more suitable rays by conducting multiple optimizations in advance. Nevertheless, we recommend exercising caution when considering PEARL-e if no a-priori assumptions about the Pareto front can be made,. and PEARL-NdS allows the user to workaround this challenge.
@@
To sum up, for the unconstrained problems with two and three objectives, the PEARL-NdS approach, in particular the crowding-based one, exacerbates high HV but also $I_{C}$, and C_metric surpassing NSGA-III systematically. PEARL-$\epsilon$ may presents high $I_{C}$, but the solutions found are usually too localized, which resulted in a lower HV. PEARL-e always under-performed, and tweaking the $\alpha$ might help improving it. However, for real-world application, a practitioner may prefer to limit the amount of tuning, especially for problems encountered for the very first time. For all of these reasons, we will pursue the study with PEARL-NdS.
@@ 5.2.2 Constrained Benchmark problems
For the constraints, we consider $F_{q}\leq 1.85$, $C_{b}\leq 1200$ ppm, and Peak Bu $\leq 62$ GWd/tHm, similar to [1, 2]. For the following with omit the ”C-” in front of ”PEARL” in the text and in the figures. The rank-based method, PEARL-NdS (niching2) and PEARL-NdS (crowding2) are using CL by default, hence we omit to mention it every time. All results with PEARL-NdS cases using CL will be showcased. Here, 50,000 samples will be collected for both the Two- and Three-obj cases. We use $\lambda$_ equal to 32, $\mu$ equal to 2, mutpb to 0.3, and cxpb to 0.65 in the both cases. We found that these hyper-parameters provided a good trade-off between exploration and exploitation for the NSGA in each instance to satisfy the constraints and then optimize in the feasible space. Since PEARL will propose more solutions from the different buffers, we will provide the pareto front of all the solutions ever found from NSGA-III rather than the last population (i.e., 32 of them) for fairness. The individual results for the Two- and Three-obj cases are given in Figure 3 on the top and bottom, respectively (grouped in five figures each). The hyper-parameters utilized are the same as the previous section. We use M equal to $\kappa$ (= 128).
@@
Figures 4 and 5 illustrate the evolution of various Figures of Merit (FOMs) of interest across 100 generations, where one generation corresponds to 500 computational samples. These figures demonstrate that, in alignment with the trade-offs observed in Figure 1 in the unconstrained case, there exists a trade-off between $L_{C}$ and $F_{\Delta h}$ for low values of the latter but it diminishes for high values. In the former case, the pattern may be asymmetric incurring high $F_{\Delta h}$ but also susceptible to suboptimal utilization of certain FAs, resulting in reduced cycle energy. The examples mentioned above provide straightforward first-order (i.e., linear) explanations. However, there are additional alternative solutions available by manipulating factors such as enrichment, burnable poison, and their placement within the core, which introduce numerous trade-offs at a more detailed level.
@@
Among the resulting best solutions, which include $[1.409,500.3]$, $[1.414,511.5]$, [1.416,512.][1.416,512.][ 1.416 , 512 . ], $[1.417,513]$, and $[1.418,513.4]$ all were found with the PEARL-NdS (crowding) with CL method exclusively. Additionally, these solutions conform to the feasibility constraints outlined in [1, 2]. In Table 10, you can find additional MOO metrics pertaining to the Two-obj case. As previously mentioned, despite the distance-based methods uncovering a smaller proportion of feasible solutions (as indicated in the last row of the Table and the top-left plot in the figure of 4), they excel in discovering a greater diversity among these solutions. Consequently, the PEARL-NdS (crowding) and PEARL-NdS (niching) approaches exhibit higher HV values. The former also boast the highest number of solutions belonging to the Pareto front (as indicated by the $I_{C}$ in the Table) although niching2 achieve the lowest IGD, GD, and $I_{\epsilon}^{+}$ values.
Figure 3: Two-obj (top) and Three-obj (bottom) solutions generated for all the algorithms utilized. The Pareto front is a line for the Two-obj case, while it is drawn with scattered points for the Three-objective one.
@@
Figure 4: Evolution of the FOMs (considered as constraints in [1, 2]) over 100 generations for the Two-obj case for PEARL-NdS (crowding) for the unconstrained, crowding,niching,niching2, and crowding2. The unconstrained case is also provided (purple curves). The lighter areas cover the average ±$\pm\sigma$.
Figure 5: Evolution of the FOMs (considered as constraints in [1, 2]) over 100 generations for the Three-obj case for PEARL-NdS (crowding) for the unconstrained, crowding,niching,niching2, and crowding2. The unconstrained case is also provided (purple curves). The lighter areas cover the average ±$\pm\sigma$.
@@
FOMs PEARL-NdS with CL crowding / niching PEARL-NdS niching2 / crowding2 NSGA-III Optimal direction
HV 2.541 / 2.4133 2.3196 / 2.301 0.948 higher
GD 0.144 / 0.0841 0.0228 / 0.334 1.664 lower
IGD 1.784 / 2.846 1.256 / 1.842 3.735 lower
$I_{\epsilon}^{+}$ 0.046 / 0.0350 0.0270 / 0.039 0.083 lower
$I_{C}$ 11 / 2 7 / 4 0 higher
C_metric 0.6875 / 0.3333 0.7 / 0.667 0 higher
Feasible 28656 / 27115 29288 / 31039 201 higher
@@
FOMs PEARL-NdS with CL crowding / niching PEARL-NdS niching2 / crowding2 NSGA-III Optimal direction
HV 2.000 / 2.134 1.347 / 2.222 0.267 higher
GD 0.0207 / 0.003 0.0 / 0.00725 0.0148 lower
IGD 1.878 / 2.499 27.425 / 8.714 12.776 lower
$I_{\epsilon}^{+}$ 0.0570 / 0.0180 0.0 / 0.0320 0.0774 lower
$I_{C}$ 171 / 330 67 / 40 0 higher
C_metric 0.584 / 0.948 1.0 / 0.952 0.0 higher
Total feasible 25662 / 26699 30934 / 31002 157 higher

### 15 Mar 2024 v3

@@ 5.2.1 Unconstrained Benchmark problems
The details of the geometry, the decision space $\Gamma$, and the feasible space $\mathbb{O}\subset\Gamma$ are given in [1, 2]. As mentioned in Section 4.2.2, we decided to minimize the peaking $F_{\Delta h}$, maximize the $L_{C}$, and minimize the average enrichment in the core. The first test will utilize the first two objectives, while the second test will consider all three. It will be intriguing to observe how the envelope coverage changes as we are adding a third type of trade-off surface. There, we ran the RL agents (similar hyper-parameters for RL as in [1, 2]) from scratch up to 50,000 and 100,000 objective function evaluations for the Two- and Three-obj cases, respectively, while using 32 cores (i.e. 32 agents) on an Intel Xeon Gold 5220R clocked at 2.20 GHz. For cosine similarity as uniformity for PEARL-e, the option normalized_obj is set to False. Similarly, we ran the NSGA-III on 32 cores, which results in 32 individuals in a population. For the GA portion of the NSGA-III algorithms, we used the Evolutionary ($\mu$,$\lambda$) Strategy (ES) [62]. We use $\lambda$_ equal to 32, $\mu$ equal to $\lambda$_,mutpb to 0.3, and cxpb to 0.65. For SA and TS we use nChain equal to ncores. Specifically for SA we have chain_size equal to 50,lmbda to 15, alpha to 2,threshold to 0.00, chi to 0.1, and Tmin to 0.005. Lastly for TS we have penalization_weight equal to 6.0, tabu_tenure to 162, chain_size to 10, and intensification to 0.2. Since PEARL will propose more solutions from the different buffers, we will provide the Pareto front of all the solutions ever found from SO. Typically, we looked at all the solutions rather than the last population for NSGA or the last element of each chain for SA and TS (i.e., 32 of them for SO) for fairness. The individual results for the Two- and Three-obj cases are given in Figure 1 (grouped in six figures each).
@@
In the Two-obj case, a trade-off between the $L_{C}$ and $F_{\Delta h}$ becomes apparent. There exists a trade-off for low values of the latter but it diminishes for high values. In the former case, the pattern may be asymmetric incurring high $F_{\Delta h}$ but also susceptible to suboptimal utilization of certain FAs, resulting in reduced cycle energy. The examples mentioned above provide straightforward first-order (i.e., linear) explanations. However, there are additional alternative solutions available by manipulating factors such as enrichment, burnable poison, and their placement within the core, which introduce numerous trade-offs at a more detailed level. Furthermore, when peaking factors are extremely low, the $L_{C}$ can undergo significant variations, even though it is feasible to achieve a comparable $L_{C}$ level for different values of $F_{\Delta h}$, ranging from approximately 1.55 to 2.2. This situation could potentially lead to substantial economic losses if the constraint on $F_{\Delta h}$ is overly stringent. The PEARL-e and PEARL-$\epsilon$ approaches fail to find low $F_{\Delta h}$ and the latter preferentially generated higher $L_{C}$ (or lower $-L_{C}$), while the former is stuck in a low quality Pareto front. Figures 2.(a) and 2.(b) shows that all algorithms will increase the $L_{cy}$ but PEARL-e and PEARL-$\epsilon$ stays with a relatively high $F_{\Delta h}$ compared to the others, while the SO-based approaches resulted in lower $F_{\Delta h}$ more rapidly than any other PEARL. These figures alone does not provide a clear indication of which approach (PEARL-NdS (crowding), or NSGA-III) is better. However, the high HV for the PEARL-NdS (crowding) in Table 14 suggests that the latter approach could provide better volume coverage, but also surpasses the NSGA-based approach on all the other indicators. Moreover, if we only look at the PEARL-NdS (crowding) versus SO (in the figure and $I_{C}$ in Table 14), we see that PEARL will propose more Pareto-optimal points. The use of multiple metrics helps therefore to assess the performance of multi-objective approaches [49, 2] and demonstrate PEARL-NdS (crowding) superiority. Lastly, the use of ${}^{+}$ on top of IGD did not provide more information, as the solution set of PEARL-NdS (crowding) comprises most of the Pareto front, hence its IGD was already largely superior to the other algorithms.
@@
More statistical weights was given to $F_{\Delta h}$. The five best Pareto-optimal points are given by $[1.42,527.9]$ found by TS, $[1.43,543.2]$ and [1.44,546.][1.44,546.][ 1.44 , 546 . ] found by NSGA-III, $[1.45,549.1]$ found by SA, and $[1.456,553.5]$ found by the PEARL-NdS (crowding)-based approach. Among the PEARL algorithms, this provides an additional reason to consider PEARL-NdS (crowding) with CL as potentially superior. Furthermore, the best candidate has a very high $L_{C}$ compared to the limit of 500 in [1, 2]. However, approaching the limit of 1.45 by turning peaking into a constraint rather than an objective could result in lower $L_{C}$ but also discard the two best PEARL-NdS (crowding) cases mentioned above. Nevertheless, in Section 5.2.2 we will investigate the impact of introducing constraints to guide the search in PEARL for this set of problems and illustrate its potential to still yield high-quality feasible solutions.
@@
For the Three-obj case, most techniques found a Pareto-optimal point (C_metric $\neq 0.0$) except PEARL-e (cos). Nevertheless again, the PEARL-NdS (crowding) approach performs better in terms of HV as seen in Table 15. The high HV for the PEARL-NdS (crowding) technique is also illustrated by Figures 2.(c) to 2.(e). There, we observe a significant variance in the objective space explored when using the latter method, while PEARL-$\epsilon$ and PEARL-e exacerbate a very low one. Although the $\sigma$ (i.e., extent of the lighter area) over time steps is large, the mean value does not vary substantially and it resulted in solutions that are concentrated around relatively high $L_{C}$, intermediate $F_{\Delta h}$ and average enrichment. The PEARL-$\epsilon$-based approach outperforms the other algorithms in terms of GD, $I_{\epsilon}^{+}$, $I_{C}$ and C_metric but exacerbates a very poor HV, IGD, and ${}^{+}$. The latter has found lots of high-quality points (high $I_{C}$ equal to 1755 which is the highest) but these points are highly localized, as corroborated by the high IGD. Additionally, The SO-based approach (here NSGA-III) only outperforms PEARL on the ${}^{+}$. This underscores again the importance of employing multiple multi-objective metrics when analyzing the performance of the algorithms [49, 2].
@@
Figure 1: Two-obj ((a) to (f)) and Three-obj ((g) to (L))) solutions generated for all the algorithms utilized. The pareto front is a line for the Two-objective case, while it is drawn with scattered points for the Three-objective one.
@@
Table 14: Different FOMs for the Two-obj case. The nadir is $[2.653,481.2]$ and the utopia is $[1.42,558.0]$. They are obtained from cases generated over the 6 optimization runs. All the indicators are calculated only with the Pareto equivalent points. The C_metric is equal to 0 when no solution belongs to the Pareto front. $I_{C}$ is the number of Pareto-optimal points for the algorithm. The best cases are highlighted in red.
FOMs PEARL-e (cos) PEARL-$\epsilon$ PEARL NdS (crowding) NSGA-III SA TS
HV 69.805 82.337 92.054 87.786 87.623 82.922
GD 4.309 0.898 0.032 0.752 1.562 1.562
IGD 11.949 4.365 1.200 2.672 2.507 4.523
${}^{+}$ 11.652 3.574 0.003 1.882 1.572 4.403
$I_{\epsilon}^{+}$ 0.325 0.238 0.0170 0.0900 0.074 0.51
$I_{C}$ 0 0 54 2 1 1
C_metric 0.0 0.0 0.964 0.250 0.167 0.100
@@
Table 15: Different FOMs for the Three-obj case. The nadir is $[3.606,4.75,451.5]$ and the utopia $[1.43,3.911,553.4]$. They are obtained from cases generated over the 6 optimization runs. All the indicators are calculated only with the Pareto equivalent points. The C_metric is equal to 0 when no solution belongs to the Pareto front. $I_{C}$ is the number of Pareto-optimal points for the algorithm. The best cases are highlighted in red.
FOMs PEARL-e (cos) PEARL-$\epsilon$ PEARL NdS (crowding) NSGA-III SA TS
HV 84.883 75.703 98.693 97.322 89.212 86.493
GD 0.129 0.00332 0.0288 0.0386 0.093 0.107
IGD 0.621 7.156 0.139 0.547 1.830 1.063
${}^{+}$ 0.129 2.045 0.077 0.016 0.401 0.356
$I_{\epsilon}^{+}$ 0.178 0.0699 0.122 0.126 0.140 0.080
$I_{C}$ 0 1755 1152 290 29 3
C_metric 0.0 0.962 0.823 0.687 0.141 0.007
@@
Additionally, it’s noteworthy that in both optimization scenarios, PEARL-e under-performs when compared to the SO-based approaches, even though it performed well in the classical benchmark cases. This difference in performance could be attributed to the arbitrary choice of $\alpha=[1.0,1.0,1.0]$. Given that we didn’t possess prior knowledge of the Pareto front’s shape, we opted for a uniform generation of rays. It is conceivable that a designer could iteratively construct more suitable rays by conducting multiple optimizations in advance. Nevertheless, we recommend exercising caution when considering PEARL-e if no a-priori assumptions about the Pareto front can be made, and PEARL-NdS allows the user to workaround this challenge.
With regards to the SO-based approaches, NSGA-III outperforms the other methods. This was observed in our parallel works with single-objective optimization [53, 54]. There, Evolutionary Strategy (ES), a GA-based approach, outperformed SA and TS. There, our problem requires global search methods to be able to find high quality solutions rapidly, but fail to improve more in the long run because the change in one solution to the next (i.e., the crossovers) were too important. By contrast, local search methods (i.e., similar to SA and TS) only perturb few entries of the input candidate slowing down the progression of the optimization, but enables long-term steady improvements. Nevertheless, the policy in RL, which controls the number of entries of the input space perturbed, manages to be both a global search at the beginning during exploration, while act as a local search in the long-term, enabling fast and constant improvement over time. This behaviors seems to be observed in the multi-objective case as well.
@@
To sum up, for the unconstrained problems with two and three objectives, the PEARL-NdS approach, in particular the crowding-based one, exacerbates high HV but also $I_{C}$, and C_metric surpassing SO systematically, while NSGA-III is the best out of the legacy approaches. PEARL-$\epsilon$ may presents high $I_{C}$, but the solutions found are usually too localized, which resulted in a lower HV. PEARL-e always under-performed, and tweaking the $\alpha$ might help improving it. However, for real-world application, a practitioner may prefer to limit the amount of tuning, especially for problems encountered for the very first time. For all of these reasons, we will pursue the study with PEARL-NdS and NSGA-III.
@@
Figure 2: Evolution of the target FOMs for the Two-objective ((a) and (b)) and the Three-objective ((c) to (e)) problem over 100 generations for PEARL. We added vertical grid lines in the Three-objective case to help with visibility and interpretation of the results: (a) and (c) $F_{\Delta h}$, (b) and (d) $L_{cy}$, and (e) Average feed enrichment. The lighter areas cover the average ±$\pm\sigma$.
@@ 5.2.2 Constrained Benchmark problems
For the constraints, we consider $F_{q}\leq 1.85$, $C_{b}\leq 1200$ ppm, and Peak Bu $\leq 62$ GWd/tHm, similar to [1, 2]. For the following with omit the ”C-” in front of ”PEARL” in the text and in the figures. The rank-based method, PEARL-NdS (niching2) and PEARL-NdS (crowding2) are using CL by default, hence we omit to mention it every time. All results with PEARL-NdS cases using CL will be showcased. Here, 50,000 samples will be collected for both the Two- and Three-obj cases. We use $\lambda$_ equal to 32, $\mu$ equal to 2, mutpb to 0.3, and cxpb to 0.65 in the both cases. We found that these hyper-parameters provided a good trade-off between exploration and exploitation for the NSGA in each instance to satisfy the constraints and then optimize in the feasible space. Since PEARL will propose more solutions from the different buffers, we will provide the pareto front of all the solutions ever found from NSGA-III rather than the last population (i.e., 32 of them) for fairness. The individual results for the Two- ((a) to (d)) and Three-obj ((e) to (i)) cases are given in Figure 3. The hyper-parameters utilized are the same as the previous section. We use M equal to $\kappa$ (= 128).
@@
Among the resulting best solutions, which include $[1.409,500.3]$, $[1.414,511.5]$, [1.416,512.][1.416,512.][ 1.416 , 512 . ], $[1.417,513]$, and $[1.418,513.4]$ all were found with the PEARL-NdS (crowding) with CL method exclusively. Additionally, these solutions conform to the feasibility constraints outlined in [1, 2]. In Table 16, you can find additional MOO metrics pertaining to the Two-obj case. As previously mentioned, despite the distance-based methods uncovering a smaller proportion of feasible solutions (as indicated in the last row of the Table), they excel in discovering a greater diversity among these solutions. Consequently, the PEARL-NdS (crowding) and PEARL-NdS (niching) approaches exhibit higher HV values. The former also boast the highest number of solutions belonging to the Pareto front (as indicated by the $I_{C}$ in the Table) although niching2 achieve the lowest IGD, GD, and $I_{\epsilon}^{+}$ values.
Figure 3: Two-obj ((a) to (d)) and Three-obj ((e) to (i)) feasible solutions generated for all the algorithms utilized. The Pareto front is a line for the Two-obj case, while it is drawn with scattered points for the Three-objective one.
@@
FOMs PEARL-NdS with CL niching / crowding PEARL-NdS niching2 / crowding2 NSGA-III Optimal direction
HV 2.4133 / 2.541 2.3196 / 2.301 0.948 higher
GD 0.0841 / 0.144 0.0228 / 0.334 1.664 lower
IGD 2.846 / 1.784 1.256 / 1.842 3.735 lower
${}^{+}$ 1.942 / 1.732 0.042 / 0.020 3.3735 lower
$I_{\epsilon}^{+}$ 0.0350 / 0.046 0.0270 / 0.039 0.083 lower
$I_{C}$ 2 / 11 7 / 4 0 higher
C_metric 0.3333 / 0.6875 0.7 / 0.667 0 higher
Feasible 27115 / 28656 29288 / 31039 201 higher
@@
FOMs PEARL-NdS with CL niching /crowding PEARL-NdS niching2 / crowding2 NSGA-III Optimal direction
HV 2.134 / 2.000 1.347 / 2.222 0.267 higher
GD 0.003 / 0.0207 0.0 / 0.00725 0.0148 lower
IGD 2.499 / 1.878 27.425 / 8.714 12.776 lower
${}^{+}$ 0.250 / 0.030 27.423 / 8.526 10.079 lower
$I_{\epsilon}^{+}$ 0.0180 / 0.0570 0.0 / 0.0320 0.0774 lower
$I_{C}$ 330 / 171 67 / 40 0 higher
C_metric 0.948 / 0.584 1.0 / 0.952 0.0 higher
Total feasible 26699 / 25662 30934 / 31002 157 higher